OpenStudy (anonymous):

Combining like terms simplify each expression: 4n-n=0

OpenStudy (anonymous):

3n=0 n=0

OpenStudy (slaaibak):

n(4-1)=0 3n=0 n=0

myininaya (myininaya):

4n-n=4n-1n=(4-1)n=3n he is simplifying expressions not solving equations hes asking if the answer is 0 and it is not answer is 3n